# Practitioner Motives to Use Different Hyperparameter Optimization Methods

### v4

@@ 6. Conclusion
This work calls for more user-centered research on HPO, particularly on exploring purposeful ways to involve practitioners in programmatic HPO methods, decision support systems for HPO, and enhancing transparency and comprehensibility of programmatic HPO methods. We will build on the findings presented in this work and seek to identify functional human heuristics (Gigerenzer and Brighton, 2009) applied in HPO. After identifying human heuristics (e.g., (Godbole et al., 2023b)), we aim to implement them in algorithms for AutoML and evaluate those algorithms in comparison to the performance of human decision-making and less human-centered HPO tools.
@@ A.2.1. Hyperparameter Optimization in a Selected Machine Learning Project
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: Question 4 must be answered ‘Yes’. Otherwise, the interviewee must not participate in the study.
@@ A.2.2. Motives to Use the Selected Hyperparameter Optimization Methods
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: This section builds on the responses from the previous section. The used HPO methods determine whether the questions need to be inverted by using ‘(not)’, as indicated in the following questions.
@@ A.4. Debriefing
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: Please summarize the key findings from the interview.
@@
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: Outline the next steps.
@@
Author Statement
This work is not under review at another journal or conference. It has not been published in conference proceedings or journal volumes. A previous version of this work was uploaded to a https://arxiv.org/abs/2203.01717.

### v5

@@ 6. Conclusion
This work calls for more user-centered research on HPO, particularly on exploring purposeful ways to involve practitioners in programmatic HPO methods, decision support systems for HPO, and enhancing transparency and comprehensibility of programmatic HPO methods. We will build on the findings presented in this work and seek to identify functional human heuristics (Gigerenzer and Brighton, 2009) applied in HPO. After identifying human heuristics (e.g., (Godbole et al., 2023a)), we aim to implement them in algorithms for AutoML and evaluate those algorithms in comparison to the performance of human decision-making and less human-centered HPO tools.
@@ A.2.1. Hyperparameter Optimization in a Selected Machine Learning Project
Interviewer Instruction: Question 4 must be answered ‘Yes’. Otherwise, the interviewee must not participate in the study.
@@ A.2.2. Motives to Use the Selected Hyperparameter Optimization Methods
Interviewer Instruction: This section builds on the responses from the previous section. The used HPO methods determine whether the questions need to be inverted by using ‘(not)’, as indicated in the following questions.
@@ A.4. Debriefing
Interviewer Instruction: Please summarize the key findings from the interview.
@@
Interviewer Instruction: Outline the next steps.